A bank teller cashed a check for $630 using $10 bills and $20 bills. The teller handed 21 more $10 bills than $20 bills to the customer, Determine how many bills of each denomination were handed to the customer.
Answer by dabanfield(803)

A bank teller cashed a check for $630 using $10 bills and $20 bills. The teller handed 21 more $10 bills than $20 bills to the customer, Determine how many bills of each denomination were handed to the customer.
Let x be the number of $10 bills. Then x - 21 = the number of $20 bills.
So we have:
10*x + 20*(x-21) = 630
10*x + 20*x-420 = 630
30*x = 1050
x = 35 (ten dollar bills)
x-21 = 14 (twenty dollar bills)